# Fluid flow channeling and mass transport with discontinuous porosity distribution

### v3

@@ 2.3. Model parameters
The model parameters can be derived by non-dimensionalizing the physical models (1) and (2) with values given in Table 1. Using $x$ ($10^{4}\,\mathrm{m}$), $\delta\!\rho g$ and $\eta_{\mathrm{b}}$ as independent scales reduces our model parameters to
$Q=\frac{1}{60},\quad\eta_{\mathrm{b}}=1,\quad\frac{k_{\mathrm{b}}}{\mu\phi_{\mathrm{b}}^{n}}=1000,\quad\delta\!\rho g\vec{e}_{d}=\begin{pmatrix}0\\ 1\end{pmatrix},\quad T=236.682.$
Furthermore, we consider $\sigma$, as suggested in [16, 15], which is an expression of the form
@@
and provides a phenomenological model for decompaction weakening. Here $c_{1}\in(0,1]$ and $c_{2}>0$, where $1+\tanh$ can be regarded as a smooth approximation of a step function taking values in the interval $(0,2)$. In the most well-studied case $c_{1}=1$, as considered in [29], one observes the formation of porosity waves, whereas the case of $c_{1}<1$ with appropriate problem parameters and initial conditions, one can observe the formation of channels. With parameters from the stated ranges, $\sigma$ as in (6) fulfills [2, Assumptions 1], and hence we refer to [2, Sec. 4] for the well-posedness of (1).
@@
$\sigma^{\mathrm{a}}(v)=1,\qquad\sigma^{\mathrm{b}}(v)=\frac{0.002+\exp(1000v)}{1+\exp(1000v)},\qquad v\in\mathbb{R}.$
@@ 3. Hydro-mechanical model results
In this part we show numerical results for solving (1) using the method described in Section 2.1. We compare three different initial porosities which are shown in Figure 2, reflecting our main focus on investigating the cases of jump discontinuities in the initial porosity distribution. Taking an initial homogeneous porosity as a reference case (Figure 2 (a)), we consider a drop (Figure 2 (b)) versus an increase (Figure 2 (c)) in the initial porosity distribution. For the effective pressure $p$, we assume homogeneous initial data $p_{0}(x)=0$.
@@
Figure 3. Porosity and effective pressure after $T=1.5\,\mathrm{Myr}$ for a smooth initial condition ($\phi_{0}^{\mathrm{a}}$) without decompaction weakening ($\sigma^{\mathrm{a}}$(a,c) and with decompaction weakening ($\sigma^{\mathrm{b}}$(b,d)
@@
Figure 4. Porosity and effective pressure after $T=1.5\,\mathrm{Myr}$ for $\phi_{0}^{\mathrm{b}}$ (a,c) and $\phi_{0}^{\mathrm{c}}$ (b,d) without decompaction weakening ($\sigma^{\mathrm{a}}$)
Figure 5. Cross section of Figure 4 for $x_{1}=5\,\mathrm{km}$
Both results consider the case without decompaction weakening. One can see very sharp transitions of $\phi$ at the locations of the initial discontinuities. As it is especially visible for the porosity distribution, the sign of its transition (positive or negative) depends on whether the initial porosity of the upper layer was smaller (negative jump) or larger (positive jump) compared to the porosity of the underlying layer. Figure 5 shows a cross section of Figure 4 that shows the discontinuities in $\phi$ more clearly. In contrast, the solution for $p$ is continuous which aligns with the theory derived in [2, Sec. 4].
@@
Figure 6. Porosity and effective pressure after $T=1.5\,\mathrm{Myr}$ for $\phi_{0}^{\mathrm{b}}$ (a,c) and $\phi_{0}^{\mathrm{c}}$ (b,d) with decompaction weakening ($\sigma^{\mathrm{b}}$)
This can be expected: we see a narrower channel in the less porous domain, but it spreads quickly once the fluid enters the domain of high porosity and permeability. This shows that the fluid does not need to channelize as much as in the less porous domain in order to travel upwards.
@@ 4. Chemical-transport model results
An extension of the general model (1) is to consider the transport of a chemical tracer by solving (2). This is achieved by following its characteristics as described in Section 2.2. The natural range of partition coefficients can be very large [10] and their magnitudes depend on several parameters [12]. Without loss of generality, we examine the ratio $K_{D}=10^{-3}$ (as already indicated in Table 1) to consider incompatible elements. Incompatible elements are those that partition preferentially in the fluid. Solving for $\mathcal{C}$ and using the prescribed values for $\rho^{\mathrm{s}}$, $\rho^{\mathrm{f}}$ and $K_{D}$ from Table 1 directly yields $\chi^{\mathrm{s}}$ and $\chi^{\mathrm{f}}$ as well.
In this part, we will only plot the normalized chemical tracer $\mathcal{C}/\mathcal{C}_{0}$. For simplicity, we consider constant initial data $\mathcal{C}_{0}(x)=1$ since $\mathcal{C}$ in (2) can be scaled arbitrarily without affecting the solution.
@@
We see the distribution of an incompatible element that prefers to stay with the fluid, and hence, it gets transported efficiently while draining the area of origin. The role of decompaction weakening becomes more apparent in the case the channelization of the fluid flow as shown in Figure 7 (b). In that case, we observe a more pronounced enrichment in the region defined by the fluid-rich channel. It is important to note that this enrichment occurs in both the solid and the fluid, and it occurs at the expense of the trace element’s distribution in the source region.
@@
Figure 8. $\mathcal{C}/\mathcal{C}_{0}$ after $T=1.5\,\mathrm{Myr}$ for $\phi_{0}^{\mathrm{b}}$ (a) and $\phi_{0}^{\mathrm{c}}$ (b) without decompaction weakening ($\sigma^{\mathrm{a}}$)
This figure is calculated based on the HM model (porosity-pressure evolution) shown in Figure 4 and assumes no decompaction weakening ($\sigma^{\mathrm{a}}$). The results generally agree with the previous findings that show that the incompatible elements ($K_{D}=10^{-3}$) travel further and enrich the upper layer. Furthermore, this enrichment seems to be traveling slightly faster in the region where the initial fluid content was higher (central region of the domain). This requires that the propagation velocity of the enrichment front is not constant and moves further from the location of the porosity discontinuity, which is located exactly at the middle of the domain ($x_{2}=10\,\mathrm{km}$). However, the enrichment is more pronounced in the case of a negative porosity jump (case $\phi_{0}^{\mathrm{b}}$) compared to the case of positive porosity jump ($\phi_{0}^{\mathrm{c}}$).
Figure 9. $\mathcal{C}/\mathcal{C}_{0}$ after $T=1.5\,\mathrm{Myr}$ for $\phi_{0}^{\mathrm{b}}$ (a) and $\phi_{0}^{\mathrm{c}}$ (b) with decompaction weakening ($\sigma^{\mathrm{b}}$)
Finally, Figure 9 shows the resulting normalized tracer element $\mathcal{C}/\mathcal{C}_{0}$ for the case of decompaction weakening ($\sigma^{\mathrm{b}}$) and an initially discontinuous porosity $\phi_{0}$. The associated HM model can be found in Figure 6. The resulting cases show marked differences and can be summarized as follows. The case with negative jump discontinuity ($\phi_{0}^{\mathrm{b}}$) shows a marked enrichment with respect to the incompatible element. This enrichment is in general heterogeneous and it propagates further by mostly following the fluid channel. Interestingly, for the case of positive jump discontinuity ($\phi_{0}^{\mathrm{c}}$), the enrichment of the incompatible element is localized close to the discontinuity location. This is explained by the fact that the fluid spreads beyond this point as it was shown in Figure 6 (b,d).
@@ 5. Discussion and Conclusions
An additional advantage of the space-time method is that the one-way coupling of the HM problem to the CT problem can be easily solved using the pre-calculated results of the HM problem. This allows for the investigation of the behavior of various trace elements and the overall mass transport in rock formations that have discontinuous porosity. Our results confirm previous data which suggest that incompatible elements are the most mobile and can travel together with the fluid [19]. This selective enrichment in incompatible elements becomes more prominent in cases where the flow is channelized, leading to the formation of localized geochemical and mineralogical anomalies. Although such mechanisms have been discussed in previous works [24, 21], the mechanism for the channeling in our case is different. In the aforementioned studies, the formation of channels was due to the selective dissolution of matrix minerals [21, 25]. In contrast, in our case the channeling is the result of decompaction weakening [5, 31, 32]. In any case, it becomes apparent that, whatever the localization mechanism may be, the localization of the fluid amplify the enrichment of incompatible elements significantly.
The consideration of our new method in cases of initially discontinuous porosity distributions provided us with new insights with respect to the transport of incompatible elements. In particular, for the case of a negative jump discontinuity ($\phi_{0}^{\mathrm{b}}$), the discontinuity creates a chemical-enrichment front that migrates away from the location of the actual discontinuity of the HM-problem solution. Finally, it is evident from our results that the effects of the channeling of the flow together with the presence of initial discontinuities will produce a variety of element-enrichment patterns that can be investigated in future studies that focus on particular element behavior.
@@
All the Julia scripts and data necessary to reproduce the results and figures of this contribution are provided on Zenodo (https://doi.org/10.5281/zenodo.13986983).

### v4

@@ 2.3. Model parameters
The model parameters can be derived by non-dimensionalizing the physical models (1) and (3) with values given in Table 1. Choosing the independent scales
$x^{\mathrm{sc}}=10^{4}\,\mathrm{m}\,,\qquad\delta\!\rho^{\mathrm{sc}}g^{\mathrm{sc}}=5\cdot 10^{3}\,\mathrm{kg}\cdot\mathrm{m}^{-2}\cdot\mathrm{s}^{-2}\,,\qquad\eta_{\mathrm{b}}^{\mathrm{sc}}=10^{19}\,\mathrm{Pa}\cdot\mathrm{s}\,$, (15)
yields the dependent scales
$\begin{split}p^{\mathrm{sc}}&=\delta\!\rho^{\mathrm{sc}}g^{\mathrm{sc}}x^{\mathrm{sc}}=5\cdot 10^{7}\,\mathrm{Pa}\,,\\ t^{\mathrm{sc}}&=\frac{\eta_{\mathrm{b}}^{\mathrm{sc}}}{p^{\mathrm{sc}}}=2\cdot 10^{11}\,\mathrm{s}\,,\\ \frac{k_{\mathrm{b}}^{\mathrm{sc}}}{\mu^{\mathrm{sc}}}&=\frac{(x^{\mathrm{sc}})^{2}}{\eta_{\mathrm{b}}^{\mathrm{sc}}}=10^{-11}\,\mathrm{m}^{2}\cdot\mathrm{Pa}^{-1}\cdot\mathrm{s}^{-1}\,.\end{split}$ (16)
Hence we end up with the nondimensional parameters
$\breve{\eta_{\mathrm{b}}}=1\,,\quad\frac{\breve{k_{\mathrm{b}}}}{\breve{\mu}\phi_{\mathrm{b}}^{n}}=1000\,,\quad\breve{\delta\!\rho}\breve{g}\vec{e}_{d}=\begin{pmatrix}0\\ 1\end{pmatrix}\,,\quad B\,\breve{Q}=Q\delta\!\rho^{\mathrm{sc}}g^{\mathrm{sc}}x^{\mathrm{sc}}=\frac{1}{60}\,$, (17)
where $B=Q^{\mathrm{sc}}\delta\!\rho^{\mathrm{sc}}g^{\mathrm{sc}}x^{\mathrm{sc}}$ denotes a non-dimensional number that is the ratio of buoyancy stress to bulk modulus. Note that this may look similar to the Deborah number defined in [7], however, the lengthscale $x^{\mathrm{sc}}$ is taken as an independent quantity in our approach, while in other studies it is derived from the compaction length  [7, 35].
Due to the given length scale and time scale, we can directly translate physical times and domain sizes into our model parameters if we divide by $x^{\mathrm{sc}}$ or $t^{\mathrm{sc}}$, respectively. For $T=1.5\,\mathrm{Myr}$, this corresponds to
$\breve{T}=T/t^{\mathrm{sc}}=\frac{1.5\cdot 10^{6}\cdot 365.25\cdot 24\cdot 60\cdot 60}{2\cdot 10^{11}}=236.682\,.$
[…]
@@
and provides a phenomenological model for decompaction weakening. Here $c_{1}\in(0,1]$ and $c_{2}>0$, where $1+\tanh$ can be regarded as a smooth approximation of a step function taking values in the interval $(0,2)$. In the most well-studied case $c_{1}=1$, as considered in [35], one observes the formation of porosity waves, whereas the case of $c_{1}<1$ with appropriate problem parameters and initial conditions, one can observe the formation of channels. With parameters from the stated ranges, $\sigma$ as in (18) fulfills [4, Assumptions 1], and hence we refer to [4, Sec. 4] for the well-posedness of (1).
@@
$\sigma^{\mathrm{a}}(v)=1,\qquad\sigma^{\mathrm{b}}(v)=\frac{0.002+\exp(1000v)}{1+\exp(1000v)},\qquad v\in\mathbb{R}.$ (19)
@@ 3. Hydro-mechanical model results
In this part we show numerical results for solving the hydro-mechanical problem (1) using the method described in Section 2.1. We compare three different initial porosities which are shown in Figure 2, reflecting our main focus on investigating the cases of jump discontinuities in the initial porosity distribution. Taking an initial homogeneous porosity as a reference case (Figure 2 (a)), we consider a drop (along $x_{2}$; Figure 2 (b)) versus an increase (along $x_{2}$; Figure 2 (c)) in the initial porosity distribution. For the effective pressure $p$, we assume homogeneous initial data $p_{0}(x)=0$. Note that all calculations in this section are carried out on larger grids, namely grids of size $30\,\mathrm{km}$ in $x_{2}$ direction, to avoid problems near the upper boundary. This, however, does not affect the computation time significantly since the additional grid cells located above $20\,\mathrm{km}$ are generally not refined because $\phi$ and $p$ are almost constant there.
@@
Figure 3. Porosity (a,b) and effective pressure  (c,d) after $T=1.5\,\mathrm{Myr}$ for a smooth initial condition ($\phi_{0}^{\mathrm{a}}$) without decompaction weakening ($\sigma^{\mathrm{a}}$(a,c) and with decompaction weakening ($\sigma^{\mathrm{b}}$(b,d)
@@
Figure 4. Porosity (a,b) and effective pressure (c,d) without decompaction weakening ($\sigma^{\mathrm{a}}$) after $T=1.5\,\mathrm{Myr}$ for $\phi_{0}^{\mathrm{b}}$ (a,c) and $\phi_{0}^{\mathrm{c}}$ (b,d)
Figure 5. Cross section of Figure 4 for $x_{1}=5\,\mathrm{km}$ with porosity (a,b) and effective pressure (c,d)
Both results consider the case without decompaction weakening. One can see very sharp transitions of $\phi$ at the locations of the initial discontinuities. Note that the discontinuity itself cannot move since the model (1) was derived under the assumption that $\mathbf{v}^{\mathrm{s}}\approx 0$ and porosity is a property of the solid. This agrees with the theoretical results shown in [4, Thm. 4.6]. As it is especially visible for the porosity distribution, the sign of its transition (positive or negative) depends on whether the initial porosity of the upper layer was smaller (negative jump) or larger (positive jump) compared to the porosity of the underlying layer. Figure 5 shows a cross section of Figure 4 that shows the discontinuities in $\phi$ more clearly. In contrast, the solution for $p$ is continuous which aligns with the theory derived in [4, Sec. 4].
@@
Figure 6. Porosity (a,b) and effective pressure (c,d) with decompaction weakening ($\sigma^{\mathrm{b}}$) after $T=1.5\,\mathrm{Myr}$ for $\phi_{0}^{\mathrm{b}}$ (a,c) and $\phi_{0}^{\mathrm{c}}$ (b,d)
This can be expected: we see a narrower channel in the domain that has smaller porosity, but the channel spreads quickly once the fluid enters the domain of high porosity and permeability. This shows that the fluid does not need to channelize as much as in the less porous domain in order to travel upwards.
@@ 4. Chemical-transport model results
An extension of the general model (1) is to consider the transport of a chemical tracer by solving (3). This is achieved by following its characteristics as described in Section 2.2. The natural range of partition coefficients can be very large [13] and their magnitudes depend on several parameters [15]. However, as it was described in Section 1.2, it is reasonable to assume that the partition coefficient can be approximated as constant given a limited range of temperatures and constant mineralogical composition (that is implicitly assumed in our models). Without loss of generality, we examine the case where $K_{D}=10^{-3}$ (as already indicated in Table 1) to consider incompatible elements. Incompatible elements are those that partition preferentially in the fluid. Solving for $\mathcal{C}$ and using the prescribed values for $\rho^{\mathrm{s}}$, $\rho^{\mathrm{f}}$ and $K_{D}$ from Table 1 directly yields $\chi^{\mathrm{s}}$ and $\chi^{\mathrm{f}}$ as well.
In this part, we will only plot the normalized chemical tracer $\mathcal{C}/\mathcal{C}_{0}$. This allows us to quantify the overall enrichment or depletion of a trace element with respect to the initial configuration. Note that since we use a characteristics-based approach for the advection of chemical elements, the chemical evolution is calculated only for the areas where the characteristics are initialized. As a result, the domain where the chemical evolution is calculated, changes in time depending on the effective velocity $\mathbf{v}^{\mathrm{e}}$. For simplicity, we consider constant initial data $\mathcal{C}_{0}(x)=1$ since $\mathcal{C}$ in (3) can be scaled arbitrarily without affecting the solution.
@@
We see the distribution of an incompatible element that prefers to stay with the fluid, and hence, it gets transported efficiently while draining the area of origin. The role of decompaction weakening becomes more apparent in the case of the channelization of the fluid flow, as shown in Figure 7 (b). In that case, we observe a more pronounced enrichment in the region defined by the fluid-rich channel. It is important to note that this enrichment occurs in both the solid and the fluid, and it occurs at the expense of the trace element’s distribution in the source region.
@@
Figure 8. $\mathcal{C}/\mathcal{C}_{0}$ without decompaction weakening ($\sigma^{\mathrm{a}}$) after $T=1.5\,\mathrm{Myr}$ for $\phi_{0}^{\mathrm{b}}$ (a) and $\phi_{0}^{\mathrm{c}}$ (b)
This figure is calculated based on the HM model (porosity-pressure evolution) shown in Figure 4 and assumes no decompaction weakening ($\sigma^{\mathrm{a}}$). The results generally agree with the previous findings that show that the incompatible elements ($K_{D}=10^{-3}$) travel further and enrich the upper layer. Furthermore, this enrichment seems to be traveling slightly faster in the region where the initial fluid content was higher (central region of the domain). This requires that the propagation velocity of the enrichment front is not constant and moves further from the location of the porosity discontinuity, which is located exactly at the middle of the domain ($x_{2}=10\,\mathrm{km}$). In contrast, for the case of the negative initial-porosity jump, the enrichment is negligible and is located in the area just above the discontinuity. A marked feature of the discontinuous models shown in Figure 8 is that, exactly at the discontinuity, we observe a significant enrichment or depletion of $\mathcal{C}$ depending whether we have a drop ($\phi_{0}^{\mathrm{b}}$ ) or an increase ($\phi_{0}^{\mathrm{c}}$ ) in the initial porosity.
Figure 9. $\mathcal{C}/\mathcal{C}_{0}$ with decompaction weakening ($\sigma^{\mathrm{b}}$) after $T=1.5\,\mathrm{Myr}$ for $\phi_{0}^{\mathrm{b}}$ (a) and $\phi_{0}^{\mathrm{c}}$ (b)
Figure 10. Cross section of Figure 9 for $x_{1}=5\,\mathrm{km}$ with $\mathcal{C}/\mathcal{C}_{0}$ on the y-axis
Finally, Figure 9 shows the resulting normalized tracer element $\mathcal{C}/\mathcal{C}_{0}$ for the case of decompaction weakening ($\sigma^{\mathrm{b}}$) and an initially discontinuous porosity $\phi_{0}$. The associated HM model can be found in Figure 6. The resulting cases show marked differences and can be summarized as follows. The case with negative jump discontinuity ($\phi_{0}^{\mathrm{b}}$) shows a marked enrichment with respect to the incompatible element. In particular, there is a marked enrichment at the discontinuity (at $x_{2}=10\,\mathrm{km}$), and within the channel in general. Interestingly, for the case of positive jump discontinuity ($\phi_{0}^{\mathrm{c}}$), the enrichment of the incompatible element is localized close to the discontinuity location (but is smaller at the discontinuity itself). This is explained by the fact that the fluid spreads beyond this point as it was shown in Figure 6 (b,d).
@@ 5. Discussion and Conclusions
An additional advantage of the space-time method is that the one-way coupling of the HM problem to the CT problem can be easily solved using the pre-calculated results of the HM problem. This allows for the investigation of the behavior of various trace elements and the overall mass transport in rock formations that have discontinuous porosity. Our results confirm previous data which suggest that incompatible elements are the most mobile and can travel together with the fluid [23]. This selective enrichment in incompatible elements becomes more prominent in cases where the flow is channelized, leading to the formation of localized geochemical and mineralogical anomalies. Although channeling mechanisms have been discussed in previous works [1, 30, 25], the mechanism for the channeling in our case is different. In the aforementioned studies, the formation of channels was due to the selective dissolution of matrix minerals [25, 31]. In contrast, in our case the channeling is the result of decompaction weakening [8, 37, 38]. In any case, it becomes apparent that, whatever the localization mechanism may be, the localization of the fluid amplify the enrichment of incompatible elements significantly. Furthermore, our new results also show the the interaction of a fluid-filled channel with a jump discontinuity in the initial porosity. This example is very relevant for the case of fluid transport across heterogeneous layers. In particular, the results show a marked enrichment of the incompatible trace elements at the initial porosity discontinuity for the cases where the initial porosity exhibits a negative jump (i.e. porosity drops sharply at the transition). In the case of a positive jump, we observe a marked depletion at exactly the same location. The results indicate that both porosity and the incompatible-element enrichment, that are associated to the discontinuity, do not move over time and remain at the same location.
Our results indicate that the effects of the channeling of the flow together with the presence of initial discontinuities will produce a variety of element-enrichment patterns that can be investigated in future studies that focus on particular element behavior. These results can be very important in targeted mineral exploration and in the understanding of ore-formation processes.
@@
All the Julia scripts and data necessary to reproduce the results and figures of this contribution are provided in the Zenodo repository [6] (https://doi.org/10.5281/zenodo.13986982).
Appendix A Derivation of the chemical model
Starting with the conservation of mass in the two phases, we get
$\partial_{t}(\phi\rho^{\mathrm{f}}\chi^{\mathrm{f}}_{i})+\mathrm{div}_{x}(\phi\rho^{\mathrm{f}}\chi^{\mathrm{f}}_{i}\mathbf{v}^{\mathrm{f}}_{i})=\Gamma^{\mathrm{f}}_{i}\,$, (20a)
$\partial_{t}((1-\phi)\rho^{\mathrm{s}}\chi^{\mathrm{s}}_{i})+\mathrm{div}_{x}((1-\phi)\rho^{\mathrm{s}}\chi^{\mathrm{s}}_{i}\mathbf{v}^{\mathrm{s}}_{i})=\Gamma^{\mathrm{s}}_{i}\,$, (20b)
for $i=1,\ldots,n$ chemical elements, where $\Gamma^{\mathrm{f}}_{i}$, $\Gamma^{\mathrm{s}}_{i}$ denote reaction terms. Note that $\sum_{i=1}^{n}\chi^{\mathrm{f}}=\sum_{i=1}^{n}\chi^{\mathrm{s}}=1$ and $\Gamma^{\mathrm{f}}_{i}+\Gamma^{\mathrm{s}}_{i}=0$ for all $i=1,\ldots,n$, since chemical elements can only be exchanged between the two phases. Defining the barycentric velocities
$\mathbf{v}^{\mathrm{f}}=\sum_{i=1}^{n}\chi^{\mathrm{f}}_{i}\mathbf{v}^{\mathrm{f}}_{i}\,,\quad\mathbf{v}^{\mathrm{s}}=\sum_{i=1}^{n}\chi^{\mathrm{s}}_{i}\mathbf{v}^{\mathrm{s}}_{i}\,$, (21)
we rewrite (20)
$\partial_{t}\big(\phi\rho^{\mathrm{f}}\chi^{\mathrm{f}}_{i}\big)+\mathrm{div}_{x}\big(\phi\rho^{\mathrm{f}}\chi^{\mathrm{f}}_{i}\mathbf{v}^{\mathrm{f}}\big)+\mathrm{div}_{x}\big(\phi\rho^{\mathrm{f}}\chi^{\mathrm{f}}_{i}(\mathbf{v}^{\mathrm{f}}_{i}-\mathbf{v}^{\mathrm{f}})\big)=\Gamma^{\mathrm{f}}_{i}\,$, (22a)
$\partial_{t}\big((1-\phi)\rho^{\mathrm{s}}\chi^{\mathrm{s}}_{i}\big)+\mathrm{div}_{x}\big((1-\phi)\rho^{\mathrm{s}}\chi^{\mathrm{s}}_{i}\mathbf{v}^{\mathrm{s}}\big)+\mathrm{div}_{x}\big((1-\phi)\rho^{\mathrm{s}}\chi^{\mathrm{s}}_{i}(\mathbf{v}^{\mathrm{s}}_{i}-\mathbf{v}^{\mathrm{s}})\big)=\Gamma^{\mathrm{s}}_{i}\,$, (22b)
and use that for trace elements the diffusion fluxes obey the Fickean limit
$\rho^{\mathrm{f}}\chi^{\mathrm{f}}_{i}(\mathbf{v}^{\mathrm{f}}_{i}-\mathbf{v}^{\mathrm{f}})$ $=-D_{i}^{\mathrm{f}}\nabla_{x}(\rho^{\mathrm{f}}\chi^{\mathrm{f}}_{i})\,$, (23a)
$\rho^{\mathrm{s}}\chi^{\mathrm{s}}_{i}(\mathbf{v}^{\mathrm{s}}_{i}-\mathbf{v}^{\mathrm{s}})$ $=-D_{i}^{\mathrm{s}}\nabla_{x}(\rho^{\mathrm{s}}\chi^{\mathrm{s}}_{i})$ (23b)
for both the fluid and solid phase. For advection dominated problems, the diffusion coefficients $D_{i}$ are very small and hence we can cancel the corresponding terms in (22). Adding the resulting equations yields
$\partial_{t}\big(\phi\rho^{\mathrm{f}}\chi^{\mathrm{f}}_{i}+(1-\phi)\rho^{\mathrm{s}}\chi^{\mathrm{s}}_{i}\big)+\mathrm{div}_{x}\big(\phi\rho^{\mathrm{f}}\chi^{\mathrm{f}}_{i}\mathbf{v}^{\mathrm{f}}+(1-\phi)\rho^{\mathrm{s}}\chi^{\mathrm{s}}_{i}\mathbf{v}^{\mathrm{s}}\big)=\Gamma^{\mathrm{f}}_{i}+\Gamma^{\mathrm{s}}_{i}=0\,.$ (24)
Next, we define the concentration ratio $K_{D}^{i}=\frac{\rho^{\mathrm{s}}\chi^{\mathrm{s}}_{i}}{\rho^{\mathrm{f}}\chi^{\mathrm{f}}_{i}}$ as well as the total concentration $\mathcal{C}_{i}=(\phi+(1-\phi)K_{D}^{i})\rho^{\mathrm{f}}\chi^{\mathrm{f}}_{i}$ which allows us to rewrite (24) further as
$\partial_{t}\mathcal{C}_{i}+\mathrm{div}_{x}(\mathcal{C}_{i}\mathbf{v}^{\mathrm{e}}_{i})=0$ (25)
where $\mathbf{v}^{\mathrm{e}}_{i}=\frac{\phi\mathbf{v}^{\mathrm{f}}+(1-\phi)\mathbf{v}^{\mathrm{s}}}{\phi+(1-\phi)K_{D}^{i}}$. Note that this equation is equivalent to (3) where we omitted the index $i$ for convenience and assumed that $\mathbf{v}^{\mathrm{s}}\approx 0$.
Appendix B Handling discontinuous velocities
In order to solve (14) for a discontinuous velocity field $\mathbf{v}^{\mathrm{e}}$, we need to ensure mass balance at the discontinuity. This is normally done via the Rankine-Hugoniot jump condition (see, for example, [2, Sec. 4.3] or [16, Sec. 11.8]), which in this case (3) leads to
$c_{+}\mathbf{v}^{\mathrm{e}}_{+}\cdot\vec{n}=c_{-}\mathbf{v}^{\mathrm{e}}_{-}\cdot\vec{n}$ (26)
where $c_{+},c_{-},\mathbf{v}^{\mathrm{e}}_{+},\mathbf{v}^{\mathrm{e}}_{-}$ denote the values of $c$ and $\mathbf{v}^{\mathrm{e}}$ on both sides of the discontinuity and $\vec{n}$ is the normal vector with respect to the discontinuity. In the test cases shown in Figures 8 and 9 we have $\vec{n}=\vec{e}_{2}$, which simplifies the numerical calculations.
In practice, when running the explicit Euler code to solve (14), we ensure that each time step does not advect the total concentration accross the discontinuity. Once the total concentration reaches the discontinuity, we recalculate the mass flux and use it to evaluate the concentration jump (without loss of generality we call it $c_{-}$), as follows
$c_{+}=c_{-}\frac{\mathbf{v}^{\mathrm{e}}_{-}\cdot\vec{n}}{\mathbf{v}^{\mathrm{e}}_{+}\cdot\vec{n}}\,.$ (27)
Appendix C Comparison with finite difference code
Here we compare our space-time approach with a classical finite difference scheme. Note that for simplicity we chose a one-dimensional test case without decompaction weakening. Hence, this is similar to the tests shown in Figure 3 (a,c) and 4.
Figure C.1. Errors of a finite difference approximation of the one-dimensional hydromechanical model without decompaction weakening for porosity (a) and effective pressure (b)
In Figure C.1 one can see the error of the finite difference scheme at the terminal time; note the very different rates for the discontinuous and continuous test cases. This difference is more pronounced for the porosity in Figure C.1 (a) since the correct solution is discontinuous whereas the corresponding effective pressure is still continuous with a kink at the location of the discontinuity. Hence, the convergence rates for the pressure in Figure C.1 (b) are higher than for the porosity, even though they are still lower than in the case of a continuous initial porosity. Note also that the rates in the continuous case are the theoretically optimal ones. In comparison, the space-time approach does not suffer from slow rates in discontinuous cases as one can see in Figure C.2. In addition, the rates here are optimal as well.
Figure C.2. Space-time errors of the one-dimensional hydromechanical model without decompaction weakening for porosity (a) and effective pressure (b)
However, we would like to emphasize that a direct comparison is not possible. This is because the space-time approach is fundamentally different from the finite difference one. For the finite differences, we measure the $L_{2}(\Omega)$-error at the terminal time, whereas for the space-time approach we considered a more complicated space-time error norm. In addition, the number of degrees of freedom (dofs) is not directly comparable since in Figure C.1 they correspond to the number of dofs at the terminal time (even though before there were many time steps involved) and in Figure C.2 they correspond to the total number of dofs for the entire space-time grid. Finally, we note that the norms measuring the errors of the porosity in Figure C.2 (a) and of the effective pressure in Figure C.2 (b) are of different kind. In summary, even though the convergence rates of the two methods are not directly comparable, the new space-time approach does not suffer from reduced convergence rates in the presence of discontinuous initial porosities.
Appendix D Continuous approximation
In this section we want to compare the approximation of the hydromechanical and chemical model for the discontinuous initial function $\phi_{0}^{\mathrm{b}}$ and a continuous approximation of it. In Figure D.1, we plot a cross section of both the discontinuous initial function and its smooth approximation.
Figure D.1. Cross section of initial porosity $\phi_{0}^{\mathrm{b}}$ (a) and its smooth approximation (b) for $x_{1}=5\,\mathrm{km}$
In Figure D.2 (a,c), one can see the hydromechanical model solution to discontinuous case (as in Figure 6 (a,c)) whereas in Figure D.2 (b,d) the results were obtained using the continuous (even though very steep) initial function shown in Figure D.1 (b).
Figure D.2. Porosity  (a,b) and effective pressure (c,d) with decompaction weakening after $T=1.5\,\mathrm{Myr}$ for $\phi_{0}^{\mathrm{b}}$ (a,c) and its continuous approximation (b,d)
Here, one can see no major difference between the two approaches. This shows that the solution of the continuous approximation indeed approximates the discontinuous one if the continuous function is steep enough.
Figure D.3. Cross section of Figure D.2 for $x_{1}=5\,\mathrm{km}$ with porosity (a,b) and effective pressure (c,d)
However, as it is shown in the cross section in Figure D.3, the continuous approach still misses most of the steep gradient of $\phi$ at the position of the discontinuity. This can be improved by using an even steeper approximation of $\phi_{0}^{\mathrm{b}}$, which, on the other side, increases the computational complexity and slows the computation times.
A very similar behavior can be observed when looking at the chemical enrichment patterns connected to the hydromechanical model results. The chemical enrichment patterns are shown in Figure D.4 and the corresponding hydromechanical models are shown in Figure D.2.
Figure D.4. $\mathcal{C}/\mathcal{C}_{0}$ with decompaction weakening ($\sigma^{\mathrm{b}}$) after $T=1.5\,\mathrm{Myr}$ for $\phi_{0}^{\mathrm{b}}$ (a) and its continuous approximation (b)
Here, the smooth approximation results in a slightly blurred version compared to the discontinuous problem. Only when looking at the cross section in Figure D.5, we see a considerable difference of the enrichment at the discontinuity (at $x_{2}=10\,\mathrm{km}$) which can, for example, have a significant impact on the creation of ore deposits at specific layers in the subsurface.
Figure D.5. Cross section of Figure D.4 for $x_{1}=5\,\mathrm{km}$ with $\mathcal{C}/\mathcal{C}_{0}$ on the y-axis. (a) corresponds to the discontinuous solution whereas (b) corresponds to the continuous approximation